# Efficient Exact Algorithms for Minimum Covering of Orthogonal Polygons with Squares

### v1

@@ 4.1 Checking if a Set of Rec-Packs covers the Input Orthogonal Polygon
Given an orthogonal polygon $P$ and a set $\mathcal{R}$ of $|\mathcal{R}|=D$ rec-packs, where =$\mathcal{R}=\{R_{1},R_{2},\ldots,R_{D}\}$, there is an algorithm that checks if $\mathcal{R}$ (equivalently $\mathcal{E}(\mathcal{R})$) completely covers $P$. The algorithm runs in $\mathcal{O}(D\cdot(D+n)^{2})$ time.
Proof 4.2.
Clearly, if $\mathcal{R}$ does not completely cover the entire interior of $P$, the uncovered interior of $P$ will be a collection of one or more orthogonal polygons themselves. Further the edges of these polygons must be either an edge of $P$ or an edge of some rec-pack in $\mathcal{R}$. Therefore the vertices must be intersections of two such edges (either of $P$ or of some rec-pack in $\mathcal{R}$). Moreover, the number of such intersections is upper-bounded by $(2D+n)^{2}$. Therefore we check for each such intersection $\alpha$, if all four blocks with $\alpha$ as a corner are covered by some rec-pack in $\mathcal{R}$, then we are done.
Thus, the following algorithm answers whether $\mathcal{R}$ completely covers the orthogonal polygon $P$ in $\mathcal{O}(D\cdot(D+n)^{2})$ time.
Algorithm 1 Complete Coverage Checker     Input: Orthogonal Polygon $P$, $\mathcal{R}$
for all x-coordinates $x_{0}$ of vertical edges of $P$ and $\mathcal{R}$ do $\triangleright$ $\mathcal{O}(D+n)$ loop
     for all y-coordinates $y_{0}$ of horizontal edges of $P$ and $\mathcal{R}$ do $\triangleright$ $\mathcal{O}(D+n)$ loop
         for all four blocks $p$ with $(x_{0},y_{0})$ as corner do $\triangleright$ $\mathcal{O}(1)$ loop
              if $p$ lies outside $P$ then
                  continue to next iteration
@@
              flag $\leftarrow$ FALSE $\triangleright$ To keep track if $s$ is covered
              for all $R\in\mathcal{R}$ do $\triangleright$ $\mathcal{O}(D)$ loop
                  if $p$ lies inside $R$ then $\triangleright$ $\mathcal{O}(1)$ check
                       flag $\leftarrow$ TRUE
                       break out from the current loop
@@
              if flag = FALSE then $\triangleright$ incomplete covering
                  return INCOMPLETE COVERING
              end if
         end for
     end for
end for
return COMPLETE COVERING
4.2 Detecting Unambiguous Squares
We now discuss an algorithm that takes in a square $S$, a partial solution $\mathcal{R}$ for OPCS and the input orthogonal polygon $P$ and decides if $S$ is an unambiguous square given $\mathcal{R}$. We crucially use Lemma 3.16 for the correctness of the algorithm.
Lemma 4.3.
Given an orthogonal polygon $P$, a partial solution $\mathcal{R}$ for OPCS and valid square $S$, there is an algorithm that checks if $S$ is an unambiguous square given $\mathcal{R}$ in $\mathcal{O}((n^{2}+|\mathcal{R}|n)^{3})$ time.
Proof 4.4.
We start by constructing the sets $D_{x}$ and $D_{y}$ as in Lemma 3.16. By definition, $|D_{x}|,|D_{y}|=\mathcal{O}((n+|\mathcal{R}|+1)\cdot n)=\mathcal{O}(n^{2}+|\mathcal{R}|n)$. Definition 3.10 implies that $S$ is an unambiguous square given $\mathcal{R}$ if and only if there is a simplicial node $p$ in $G^{\mathcal{R}}(P)$ such that $S$ covers $p$ and all its neighbours in $G^{\mathcal{R}}(P)$. From Lemma 3.16, this is not the case if and only if all maximal squares with a corner in $D_{x}\cup D_{y}$ which cover some block not covered by $\mathcal{R}\cup\{S\}$, together cover up the entire region inside $S$ not covered by $\mathcal{R}$. In other words, all blocks $p_{1}$ in $S$ not covered by $\mathcal{R}$ have an adjacent block $p_{2}$ in $G^{\mathcal{R}}(P)$ and therefore $p_{1}$ cannot be a simplicial node in $G^{\mathcal{R}}(P)$ with $S$ covering all its neighbours in $G^{\mathcal{R}}(P)$. Let $\mathcal{D}$ be the set of maximal squares with a corner in $D_{x}\cup D_{y}$ that cover some node in $G^{\mathcal{R}}(P)$ which is not covered by $S$ (i.e. we do not consider those maximal squares which lie completely outside the region covered by $\mathcal{R}\cup\{S\}$). By construction, $|\mathcal{D}|\leq 4|D_{x}|+4|D_{y}|=\mathcal{O}(n^{2}+|\mathcal{R}|n)$.
There is some block $p$ in $S$, which is not covered by any rec-pack in $\mathcal{D}\cup\mathcal{R}$, if and only if $p$ is a simplicial node in $G^{\mathcal{R}}(P)$ with $S$ covering $p$ and all its neighbours, i.e. $S$ is an unambiguous square given $\mathcal{R}$ (Lemma 3.16). However, we cannot check for simpliciality in all blocks in $S$, as this can potentially be exponential. Nevertheless, let us draw the infinite horizontal and vertical lines from all corners of rec-packs of $\mathcal{R}\cup\{S\}\cup\mathcal{D}$ and from the vertices of $P$. This would form a grid like structure with $\mathcal{O}((n^{2}+|\mathcal{R}|n)^{2})$ rectangular regions. Clearly, any two blocks in the same rectangular region would be covered by the same subset of rec-packs in $\mathcal{R}\cup\{S\}\cup\mathcal{D}$. Therefore it would be sufficient to check for simpliciality in a single block in each rectangular region (we check all the corner blocks of the rectangular regions i.e. blocks having a corner which shares an x coordinate of a vertex in $P$ or a corner in $\mathcal{R}\cup\{S\}\cup\mathcal{D}$ as well as the y coordinate). Given the description above, checking for simpliciality will take $\mathcal{O}(n^{2}+|\mathcal{R}|n)$ time.
Therefore, given $\mathcal{R}$ the following algorithm answers whether $S$ is an unambiguous square in $\mathcal{O}((n^{2}+|\mathcal{R}|n)^{3})$ time.
Algorithm 2 Is Unambiguous Square     Input: $P$, $\mathcal{R}$, $S$
If $S$ is not maximal return NOT UNAMBIGUOUS $\triangleright$ $\mathcal{O}(n)$ time
Construct $D_{x}$, $D_{y}$ as in Lemma 3.16 $\triangleright$ $|D_{x}|,|D_{y}|=\mathcal{O}(n^{2}+|\mathcal{R}|n)$
$\mathcal{D}^{\prime}\leftarrow$maximal squares with a corner in $D_{x}\cup D_{y}$} $\triangleright$ $\mathcal{O}((n^{2}+|\mathcal{R}|n)\cdot n)$ time
$\mathcal{D}\leftarrow$squares in $\mathcal{D}^{\prime}$ covering a block not covered by $\mathcal{R}\cup\{S\}$} $\triangleright$ $|\mathcal{D}|=\mathcal{O}(n^{2}+|\mathcal{R}|n)$
$T\leftarrow\{(x,y)|x\text{ and }y\text{ are coordinates of corners in }P\text{or }\mathcal{R}\cup\mathcal{D}\cup\{S\}\}$
for $(x,y)\in T$ do $\triangleright$ $\mathcal{O}(|T|)=\mathcal{O}((|\mathcal{D}|+|\mathcal{R}|+n)^{2})=\mathcal{O}((n^{2}+|\mathcal{R}|n)^{2})$ loop
     for blocks $p$ with corner $(x,y)$ do $\triangleright$ at most $4$ possibilities
         if $p\in S$ and $\forall R\in\mathcal{R}\cup\mathcal{D},p\notin R$ then $\triangleright$ $\mathcal{O}(|\mathcal{D}|+|\mathcal{R}|)=\mathcal{O}(n^{2}+|\mathcal{R}|n)$ check
              return UNAMBIGUOUS $\triangleright$ $p$ is simplicial in $G^{\mathcal{R}}(P)$
         end if
     end for
end for
return NOT UNAMBIGUOUS
4.3 Generating Rec-Packs within a Strip
We wish to use Lemma 3.21 to extend an existing partial solution by adding a rec-pack. In order to do this, we need to construct rec-packs $R$ within a strip (between $e_{1}$, $e_{2}$, Lemma 3.21) which shares its shorter side (perpendicular to $e_{1}$ and $e_{2}$) with the side of a given square $S$ (we call this the ‘seed’). We discuss the following algorithm to construct rec-packs $R$ for the setting in Lemma 3.21.
Lemma 4.5.
Given a maximal square $S$ (referred to as the ‘seed’) of an orthogonal polygon $P$ with side length $d$, there is an algorithm to test if there are horizontal polygon edges $e_{1},e_{2}$ and some natural number $\eta\in\mathbb{N}$, such that
  • The top and bottom edges of $S$ overlaps with horizontal polygon edges $e_{1}$ and $e_{2}$ respectively.
  • $e_{1}$ contains the top right corner of $S$ and the distance between the top right corner of $S$ and the right endpoint $e_{1}$ is more than $\eta d$.
  • $e_{2}$ contains the bottom right corner of $S$ and the distance between the bottom right corner of $S$ and the right endpoint of $e_{2}$ is more than $\eta d$.
  • There is a strip between $e_{1},e_{2}$, and the right side of the strip is more than $\eta d$ distance away from the right side of $S$.
Moreover, if this exists, the algorithm finds the largest possible $\eta$ and if $\eta\geq 2$ the algorithm returns the rec-pack $R$ of width $d$ and strength $\eta-1$, such that the vertical sides of $R$ are of length $d$ and the right side of the square $S$ coincides with the left side of $R$. The algorithm runs in total $\mathcal{O}(n)$ time.
Note that this can be done for all four directions. Therefore, henceforth if we need to apply a similar algorithm in a different direction, we shall still refer to this Lemma.
This essentially tests if Lemma 3.21 can be applied to the setting or not; if yes, this also outputs the rec-pack $R$ with the second largest strength $\eta-1$ for which the setting applies.
Proof 4.6.
Consider the following algorithm.
Algorithm 3 Generate-Rec-pack-left     Input: $P$, seed $S$
Try finding $e_{1}$, $e_{2}$ containing top-left and bottom-right corner of $S$triangleright $\mathcal{O}(n)$ check
if $e_{1}$ or $e_{2}$ not found then return NOT-APPLICABLE end if
$x_{1}\leftarrow$ x coordinate of right side of $S$
$x_{2}\leftarrow\min(\text{x coordinate of right end of }e_{1},\text{x coordinate of right end of }e_{2})$
$x_{2}\leftarrow\min(x_{2},\text{ x coordinate $x^{\prime}$ of polygon vertex lying between }e_{1},e_{2}:x^{\prime}>x_{1}$)
$d\leftarrow$ side length of $S$
if $x_{2}-x_{1}\leq d$ then return NOT-APPLICABLE end if $\triangleright$ Setting not applicable
$\eta\leftarrow\lceil\frac{x_{2}-x_{1}}{d}\rceil-1$triangleright gets maximum possible strength $\eta$
if $\eta=1$ then return SHORT-STRIP end if
This algorithm first checks for the existence of polygonal edges $e_{1}$ and $e_{2}$ ($\mathcal{O}(n)$ loop on all polygon-edge) and further checks if $e_{1}$ and $e_{2}$ have right end points lying farther than $\eta d$ distance for at least $\eta=1$ (the minimum value). If any of these do not happen, no such $R$ exists and the algorithm outputs ‘NOT-APPLICABLE’.
Moreover, if it is applicable, it finds the largest value of $\eta$ such that $e_{1},e_{2}$ have both of their right endpoints more than $\eta d$ distance away from the right side of $S$. Finally, the algorithm returns the rec-pack $R$ of strength $\eta-1$ as needed if $\eta\geq 2$ or returns ‘SHORT-STRIP’ if $\eta=1$.
The entire algorithm runs in linear time and can be extended to all directions.
{observation}
The rec-packs generated in Algorithm 3 leaves at least distance $d$ and at most distance $2d$ from the right side of the strip contained in $e_{1}$, $e_{2}$ and can not intersect with any valid square that covers some region outside the strip containing $e_{1},e_{2}$.
The idea for generating $R$ with strength $\eta-1$ (or returning ‘SHORT-STRIP’ if $\eta=1$) is to ensure that $R$ does not intersect with any square that partially lies outside the strip in between $e_{1}$ and $e_{2}$. Moreover, if we get ‘SHORT-STRIP’ as an output, we know that there is no efficient rec-pack representation necessary as the number of squares needed to cover any such region with $\eta=1$ is at most $2$.
While using this as a subroutine for our final algorithm discussed later, we need to check that given a partial solution $\mathcal{R}\cup\{S\}$ with $S$ acting as a seed, if the generated rec-pack $R$ intersects with any rec-pack in $\mathcal{R}$. If there is no intersection, we can say that $\mathcal{R}\cup\{R\}$ is also a partial solution.
4.4 Polynomial Time Algorithm: Building up Partial Solutions
Finally, we are ready to state the polynomial time algorithm for OPCS when the input polygon is in terms of its $n$ vertices. We use the following framework: Start with an empty partial solution, keep extending the partial solution either by adding unambiguous squares (Lemma 3.12) or by adding rec-packs generated by a seed square (Algorithm 3 and Lemma 3.21). We stop this whenever we have a set of rec-packs which completely covers the polygon (detected using Algorithm 1). This will be guaranteed to be a minimum cover, because throughout the algorithm the set of rec-packs is guaranteed to be a partial solution by Lemma 3.12 and Lemma 3.21.
To find unambiguous squares given a partial solution $\mathcal{R}$ of OPCS, we generate all possible maximal squares having end points in $C_{x}\cup C_{y}$ as defined in Lemma 3.14. We are guaranteed to get an unambiguous square given $\mathcal{R}$ (due to Lemma 3.14). Next, we test for all such generated squares if it is unambiguous using Algorithm 2.
Formally, the algorithm is as follows.
Algorithm 4 Extend-Partial-Solution     Input: $P$
$\mathcal{R}\leftarrow\emptyset$triangleright Start with empty partial solution
while $\mathcal{R}$ does not cover $P$ do $\triangleright$ $\mathcal{O}(|\mathcal{R}|\cdot(|\mathcal{R}|+n)^{2})$ time check with Algorithm 1
     Construct $C_{x},C_{y}$ as Lemma 3.14 $\triangleright$ $|C_{x}|,|C_{y}|=\mathcal{O}(n\cdot(n+|\mathcal{R}|))$
     $\mathcal{S}^{\prime}\leftarrow\{\text{squares with corner in }C_{x}\cup C_{y}\}$triangleright $|\mathcal{S}^{\prime}|=\mathcal{O}(n\cdot(n+|\mathcal{R}|))$, time $=\mathcal{O}(|\mathcal{S}^{\prime}|\cdot n)$
     for $S\in\mathcal{S}^{\prime}$ do $\triangleright$ $|\mathcal{S}^{\prime}|=\mathcal{O}(n\cdot(n+|\mathcal{R}|))$ loop
         if $S$ is unambiguous in $P$ given $\mathcal{R}$ $\triangleright$ $\mathcal{O}((n^{2}+|\mathcal{R}|n)^{3})$ using Algorithm 2
              $\mathcal{R}\leftarrow\mathcal{R}\cup\{S\}$triangleright $\mathcal{R}$ is still a partial solution by Lemma 3.12
              $R_{l}\leftarrow$ generated rec-pack with seed $S$ to the left. $\triangleright$ $\mathcal{O}(n)$ with Algorithm 3
              $R_{r}\leftarrow$ generated rec-pack with seed $S$ to the right. $\triangleright$ Similar to Algorithm 3
              $R_{t}\leftarrow$ generated rec-pack with seed $S$ to the top. $\triangleright$ Similar to Algorithm 3
              $R_{b}\leftarrow$ generated rec-pack with seed $S$ to the bottom. $\triangleright$ Similar to Algorithm 3
              for $R\in\{R_{l},R_{r},R_{t},R_{b}\}$ do $\triangleright$ $\mathcal{O}(1)$ loop
                  if $R\notin$NOT-APPLICABLE, SHORT-STRIP} then $\triangleright$ Algorithm 3 output
                       if $R$ does not intersects with any rec-pack in $\mathcal{R}$ then $\triangleright$ $\mathcal{O}(n)$ check
                           $\mathcal{R}\leftarrow\mathcal{R}\cup\{R\}$triangleright $\mathcal{R}$ remains a partial solution by Lemma 3.21
                       end if
                  end if
              end for
              break out of for loop, continue next while loop iteration
@@ 4.4 Polynomial Time Algorithm: Building up Partial Solutions
$s\leftarrow 0$
for $R\in\mathcal{R}$ do $\triangleright$ $\mathcal{O}(|\mathcal{R}|)$ loop
     $s\leftarrow s+(\text{strength of }R)$triangleright $|\mathcal{E}(R)|=$ strength of $R$.
end for
return $s$
First, as the generated rec-packs are only added to $\mathcal{R}$ if they do not intersect with any rec-pack in $\mathcal{R}$, we make the following observation.
{observation}
The only non-trivial rec-packs in $\mathcal{R}$ are the ones generated by a seed, and no two non-trivial rec-packs in $\mathcal{R}$ intersect with each other.
@@
Proof 4.8.
Termination. Each iteration of the while loop takes finite time and every other step outside the while loop takes finite time as well. Moreover, the number of while loop iterations is also finite as at each step, as in each step the algorithm is guaranteed to find an unambiguous square (Lemma 3.14) and therefore the size of the partial solution increases at least by one. Further, this unambiguous square added covers at least one currently uncovered block of the polygon, thereby decreasing in the current step the area of the uncovered region of the polygon at least by $1$. Since $P$ has a finite area, the number of while loop iterations is finite, and hence Algorithm 4 terminates in finite time.
Minimum Covering. By Lemma 3.12 and Lemma 3.21, $\mathcal{R}$ is always a partial solution, and therefore if $\mathcal{R}$ covers $P$ completely, it has to be a minimum cover. This means $\mathcal{E}(\mathcal{R})$ is a minimum cover. Therefore, we have,
$\text{{OPCS}}(P)=|\mathcal{E}(\mathcal{R})|=|\bigcup\limits_{R\in\mathcal{R}}\mathcal{E}(R)|=\sum\limits_{R\in\mathcal{R}}|\mathcal{E}(R)|=\sum\limits_{R\in\mathcal{R}}(\text{strength of }R)$
which completes the proof.
Our next step is to bound the time for each iteration of the while loop.
@@
Each iteration of the while loop of Algorithm 4 runs in $\mathcal{O}((n^{2}+|\mathcal{R}|n)^{4})$ time.
Proof 4.10.
We analyse the time taken at each step.
  • Checking the condition of the while loop takes $\mathcal{O}(|\mathcal{R}|\cdot(|\mathcal{R}|+n)^{2})$ time, which is clearly bounded by $\mathcal{O}((n^{2}+|\mathcal{R}|n)^{4})$.
  • Generating $\mathcal{S}^{\prime}$, and checking for all possible squares in $\mathcal{S}^{\prime}$ if they are ambiguous takes time $\mathcal{O}((n^{2}+|\mathcal{R}|n)\cdot n+(n^{2}+|\mathcal{R}|n)\cdot(n^{2}+|\mathcal{R}|n)^{3})=\mathcal{O}((n^{2}+|\mathcal{R}|n)^{4})$ as $|S^{\prime}|=\mathcal{O}(n^{2}+|\mathcal{R}|n)$
  • Generating rec-packs with the seed $S$ and the respective checks are only done four times per while loop iteration, and therefore they take up a total overhead of $\mathcal{O}(n)$.
Hence, each iteration of the while loop in Algorithm 4 takes $\mathcal{O}((n^{2}+|\mathcal{R}|n)^{4})$ time.
We now bound the total time in terms of the size of final set of rec-packs $\mathcal{R}$.

### v2

@@ 5.1 Checking if a set of rec-packs covers an orthogonal polygon
There exists an algorithm to check in $\mathcal{O}(|\mathscr{R}|\log|\mathscr{R}|)$, whether the set of rectangles $\mathscr{R}$ completely cover an orthogonal polygon $P$.
We assumed that the area of the polygon $P$ is already known. This is a natural assumption as this can be computed in $\mathcal{O}(n)$ time using the shoelace formula or the surveyor’s area formula [9], which is asymptotically the same as just reading the polygon.
5.2 Detecting unambiguous squares
Next, we discuss an algorithm that takes in a square $S$, a partial solution $\mathscr{R}$ for OPCS and the input orthogonal polygon $P$; and decides if $S$ is an unambiguous square given $\mathscr{R}$. We crucially use Lemma 4.17 for the correctness of the algorithm.
Lemma 5.2.
Given an orthogonal polygon $P$, a partial solution $\mathscr{R}$ for OPCS and valid square $S$, there is an algorithm that checks if $S$ is an unambiguous square given $\mathscr{R}$ in $\mathcal{O}(n(n+|\mathscr{R}|)^{2})$ time.
Proof 5.3.
We assume $S$ is maximal. We start by constructing the sets $D_{x}$ and $D_{y}$ as in Lemma 4.17. By definition, $|D_{x}|,|D_{y}|=\mathcal{O}((n+|\mathscr{R}|+1)\cdot n)=\mathcal{O}(n^{2}+|\mathscr{R}|n)$. Let $\mathcal{D}$ be the set of maximal squares with a corner in $D_{x}\cup D_{y}$. Among these, let $\mathcal{D}^{\prime}\subseteq\mathcal{D}$ be the set of maximal squares which cover at least one block that is not covered by $\mathscr{R}$ or $S$.
Suppose there is a block $q$ covered by $S$, which has a neighbour $q^{\prime}$ in $G^{\mathscr{R}}(P)$ lying outside $S$. By Lemma 4.17, there will be a square $S_{q}\in\mathcal{D}^{\prime}$ which covers $q$.
$S$ is an unambiguous square given $\mathscr{R}$ if and only if there is a simplicial node $p$ in $G^{\mathscr{R}}(P)$ such that $S$ covers $p$ and all its neighbours in $G^{\mathscr{R}}(P)$ (Definition 4.9). Therefore, there is no square $S_{p}\in\mathcal{D}^{\prime}$ that covers $p$. This means that $S$ is not unambiguous if and only if the rectangles in $\mathscr{R}\cup\mathcal{D}^{\prime}$ cover up $S$ entirely (along with possibly some portion outside $S$ as well).
We can use Lemma 5.1 to actually check this in time $\mathcal{O}((|\mathscr{R}|+|\mathcal{D}^{\prime}|)\log(|\mathscr{R}|+|\mathcal{D}^{\prime}|))=\mathcal{O}((n^{2}+|\mathscr{R}|n)\log(n^{2}+|\mathscr{R}|n))$ by checking if the rectangles $\{R\cap S\mid R\in\mathscr{R}\cup\mathcal{D}^{\prime}\}$ cover $S$ completely or not.
We have $|\mathcal{D}^{\prime}|\leq|\mathcal{D}|\leq 4|D_{x}|+4|D_{y}|=\mathcal{O}(n^{2}+|\mathscr{R}|n)$. Computing $D_{x},D_{y}$ are trivial from their definitions. $\mathcal{D}$ can be computed in $\mathcal{O}(|\mathcal{D}|n)=\mathcal{O}((n^{2}+|\mathscr{R}|n)n)$ time using Lemma 4.1. Moreover, for every square in $\mathcal{D}$, it can be checked in $\mathcal{O}(|\mathscr{R}|)$ time if it lies completely inside the region defined by $\mathscr{R}\cup\{S\}$. We throw out all such squares, and just keep the rest in $\mathcal{D}^{\prime}$. Hence $\mathcal{D}^{\prime}$ can be computed in $\mathcal{O}(|\mathcal{D}^{\prime}||\mathscr{R}|)=\mathcal{O}((n^{2}+|\mathscr{R}|n)|\mathscr{R}|)$ time.
This gives us a total time complexity of $\mathcal{O}(n(n+|\mathscr{R}|)^{2})$. Putting everything together, the algorithm we get is as follows.
Algorithm 1 Check-if-Unambiguous     Input: $P$, $\mathscr{R}$, $S$
if $S$ is not maximal then $\triangleright$ $\mathcal{O}(n)$ time
     return NOT-UNAMBIGUOUS
@@ 5.2 Detecting unambiguous squares
Construct $D_{x}$, $D_{y}$ as in Lemma 4.17
$\mathcal{D}\leftarrow$maximal squares with a corner in $D_{x}\cup D_{y}$} $\triangleright$ $\mathcal{O}((n^{2}+|\mathscr{R}|n)n)$
$\mathcal{D}^{\prime}\leftarrow$squares in $\mathcal{D}$ covering a block not covered by $\mathscr{R}\cup\{S\}$} $\triangleright$ $\mathcal{O}((n^{2}+|\mathscr{R}|n)|\mathscr{R}|)$
if $\{R\cap S\mid R\in\mathscr{R}\cup\mathcal{D}^{\prime}\}$ cover $S$ completely then $\triangleright$ $\mathcal{O}((|\mathscr{R}|+|\mathcal{D}^{\prime}|)\log(|\mathscr{R}|+|\mathcal{D}^{\prime}|))$
     return NOT-UNAMBIGUOUS
else
     return UNAMBIGUOUS
end if
5.3 Generating rec-packs within a strip
Our next step would be to algorithmically use Lemma 4.22 to extend an existing partial solution by adding a rec-pack. In order to do this, we look into a maximal square $S$, which we call the ‘seed’, and check if there is a corresponding strip $Y$ (as defined in Lemma 4.22). If they exist, we construct the corresponding rec-pack which does not intersect with the current partial solution, and add it to our partial solution. The resultant set of rec-packs still form a partial solution due to Lemma 4.22. We formally discuss this algorithmic result.
Lemma 5.4.
Given a maximal square $S$ of side length $d$, referred to as the ‘seed’, and a partial solution $\mathscr{R}$ containing $S$, there is an algorithm to test if there are horizontal polygon edges $e_{1},e_{2}$ such that
  • the top and bottom sides of $S$ overlaps with horizontal polygon edges $e_{1}$ and $e_{2}$ respectively.
  • $e_{1}$ contains the top right corner of $S$
  • $e_{2}$ contains the bottom right corner of $S$
  • there is a strip $Y$ between $e_{1},e_{2}$, and the right side of the strip is more than $d$ distance away from the right side of $S$.
Moreover, if this exists, the algorithm finds and returns the rec-pack $R$ of width $d$, and the largest possible strength $\eta\geq 1$, such that:
  • the vertical sides of $R$ are of length $d$
  • right side of the square $S$ coincides with the left side of $R$
  • $R$ does not overlap with partial solution $\mathscr{R}$
The algorithm runs in total $\mathcal{O}(n+|\mathscr{R}|)$ time.
Proof 5.5.
Consider the following algorithm.
Algorithm 2 Generate-Rec-pack-left     Input: $P$, seed $S$, set of rec-packs $\mathscr{R}$
Find $e_{1}$, $e_{2}$ containing top-right and bottom-right corner of $S$triangleright $\mathcal{O}(n)$ check
if $e_{1}$ or $e_{2}$ is not found then
     return NONE
end if
$x_{1}\leftarrow$ x coordinate of right side of $S$
$x_{2}\leftarrow\min\{\text{x coordinate $x^{\prime}$ of vertex in $P$ lying between }e_{1},e_{2}\text{ such that }x^{\prime}>x_{1}$}
$x_{3}\leftarrow\min\{\text{x coordinate $x^{\prime}$ of corner in $\mathscr{R}$ lying between }e_{1},e_{2}\text{ such that }x^{\prime}>x_{1}$}
$d\leftarrow$ side length of $S$
if $\min(x_{2},x_{3})-x_{1}\leq d$ $\triangleright$ rec-pack $R$ does not exist
     return NONE
end if
$\eta\leftarrow\left\lfloor\dfrac{\min(x_{2},x_{3})-x_{1}}{d}\right\rfloor$triangleright gets maximum possible strength $\eta$
Figure 9: Setting and proof of Lemma 5.4
This algorithm first checks for the existence of polygonal edges $e_{1}$ and $e_{2}$ by looping over all polygon-edges. Next it checks if there is a valid rec-pack $R$, by checking the x-coordinate $x_{2}$ of the polygon vertex that bounds the right side of $Y$, as well as the x-coordinate $x_{3}$, of some rec-pack in $\mathscr{R}$ that restricts the right side of $R$, as they cannot overlap (Figure 9). If either of these fail, there is no rec-pack $R$, and the algorithm returns ‘NONE’.
Otherwise, it finds the largest value of $\eta$ and returns the rec-pack $R$ as required. The entire algorithm runs in time $\mathcal{O}(n+|\mathscr{R}|)$ time, as the algorithm just consists of loops over the vertices of $P$ and corners of rec-packs in $\mathscr{R}$.
Remark 5.6.
Lemma 5.4 is applicable for all four directions.
5.4 Polynomial-time algorithm: building up partial solutions
Finally, we are ready to state the polynomial time algorithm for OPCS when the input orthogonal polygon $P$ is given in terms of its $n$ vertices. We use the following framework:
  1. 1.
    Start with an empty partial solution.
  2. 2.
    Keep extending the partial solution either by adding unambiguous squares (Algorithm 1) or by adding rec-packs generated by a seed (Algorithm 2).
  3. 3.
    Terminate when $P$ is completely covered (Lemma 5.1).
This must be a minimum cover as throughout the algorithm, the set of rec-packs is guaranteed to be a partial solution (Lemma 4.11 and Lemma 4.22).
To find unambiguous squares given a partial solution $\mathscr{R}$ of OPCS, we generate all possible maximal squares having end points in $C_{x}\cup C_{y}$ as defined in Lemma 4.13. We are guaranteed to get an unambiguous square given $\mathscr{R}$ (Lemma 4.13). Next, we test for all such generated squares if it is unambiguous using Algorithm 1.
Formally, the algorithm is as follows.
Algorithm 3 Extend-Partial-Solution     Input: Orthogonal Polygon $P$ with $n$ vertices
$\mathscr{R}\leftarrow\emptyset$triangleright Empty partial solution
while $\mathscr{R}$ does not cover $P$ do $\triangleright$ $\mathcal{O}(|\mathscr{R}|\log|\mathscr{R}|)$ check, Lemma 5.1
     Construct $C_{x},C_{y}$ as Lemma 4.13 $\triangleright$ $|C_{x}|,|C_{y}|=\mathcal{O}(n(n+|\mathscr{R}|))$
     $\mathscr{S}^{\prime}\leftarrow\{\text{squares with corner in }C_{x}\cup C_{y}\}$triangleright $|\mathscr{S}^{\prime}|=\mathcal{O}(n(n+|\mathscr{R}|))$, time $=\mathcal{O}(|\mathscr{S}^{\prime}|\cdot n)$
     for $S\in\mathscr{S}^{\prime}$ do $\triangleright$ $|\mathscr{S}^{\prime}|=\mathcal{O}(n\cdot(n+|\mathscr{R}|))$ loop
         if $S$ is unambiguous in $P$ given $\mathscr{R}$ $\triangleright$ $\mathcal{O}(n(n+|\mathscr{R}|)^{2})$, Algorithm 1
              $\mathscr{R}\leftarrow\mathscr{R}\cup\{S\}$triangleright $\mathscr{R}$ remains a partial solution, Lemma 4.11
              $R_{l}\leftarrow$ generated rec-pack with seed $S$ to the left. $\triangleright$ $\mathcal{O}(n)$, Algorithm 2
              $R_{r}\leftarrow$ generated rec-pack with seed $S$ to the right. $\triangleright$ $\mathcal{O}(n)$, Algorithm 2
              $R_{t}\leftarrow$ generated rec-pack with seed $S$ to the top. $\triangleright$ $\mathcal{O}(n)$, Algorithm 2
              $R_{b}\leftarrow$ generated rec-pack with seed $S$ to the bottom. $\triangleright$ $\mathcal{O}(n)$, Algorithm 2
              for $R\in\{R_{l},R_{r},R_{t},R_{b}\}$ do
                  if $R\neq$ NONE then $\triangleright$ Algorithm 2 output
                       $\mathscr{R}\leftarrow\mathscr{R}\cup\{R\}$triangleright $\mathscr{R}$ remains a partial solution, Lemma 4.22
@@ 5.4 Polynomial-time algorithm: building up partial solutions
              break out of the for loop, continue to the next while loop iteration
@@
return $\sum\limits_{R\in\mathscr{R}}(\text{strength of }R)$triangleright $\mathcal{O}(|\mathscr{R}|)$ loop
@@
Proof 5.8.
Termination. Each iteration of the while loop takes finite time and all steps outside the while loop take finite time as well. Moreover, the number of while loop iterations is also finite, as in each step the algorithm covers some previously uncovered block. Hence Algorithm 3 terminates in finite time.
Minimum Covering. By Lemma 4.11 and Lemma 4.22, $\mathscr{R}$ is always a partial solution, and therefore if $\mathscr{R}$ covers $P$ completely, it has to be a minimum cover.
Our next step is to bound the running time of Algorithm 1 with respect to the final set of rec-packs $\mathscr{R}$.
@@
If the final set of rec-packs from Algorithm 3 is $\mathscr{R}$, then its total running time is $\mathcal{O}(n^{2}|\mathscr{R}|(n+|\mathscr{R}|)^{3})$.
Proof 5.10.
First we analyse the running time of each iteration of the while loop.
  • Checking the condition of the while loop takes $\mathcal{O}(|\mathscr{R}|\log|\mathscr{R}|)$ time.
  • Generating $\mathscr{S}^{\prime}$ takes time $\mathcal{O}(n|\mathscr{S}^{\prime}|)=\mathcal{O}(n^{2}(n+|\mathscr{R}|))$ time.
  • Checking for all possible squares in $\mathscr{S}^{\prime}$ if they are ambiguous takes time $\mathcal{O}(n(n+|\mathscr{R}|)\cdot n(n+|\mathscr{R}|)^{2})=\mathcal{O}(n^{2}(n+|\mathscr{R}|)^{3})$. This will asymptotically be the slowest step.
  • Generating rec-packs with the seed $S$ and the respective checks are only done four times per while loop iteration. Therefore they take up a total overhead of $\mathcal{O}(n+|\mathscr{R}|)$, per while loop iteration (Lemma 5.4).
Hence, each iteration of the while loop takes $\mathcal{O}(n^{2}(n+|\mathscr{R}|)^{3})$ time. Moreover, each iteration increases the size of $\mathscr{R}$ by at least $1$. So the total running time of Algorithm 3 should be $\mathcal{O}(|\mathscr{R}|\cdot n^{2}(n+|\mathscr{R}|)^{3})$.
The dependence of the running time on the final set of rec-packs $\mathscr{R}$, makes it seem like an output-sensitive algorithm. However, we bound the size of $\mathscr{R}$ to be quadratic in $n$.